given that a(vector)=i(cap)-j(cap) and b(vector)=i(cap)+2j(cap) are two vectors find a unit vector coplanar with a(vector) and b(vector) and perpendicular to a(vector)

Grade:12

2 Answers

Sher Mohammad IIT Delhi
10 years ago
a=i-j
b=i+2j
c=xi+yj
a.c=x-y=0;
x=y
c=xi+xj
the unit vector is 1/sqrt(2)(i+j)

thanks
sher mohammad
b.tech, iit delhi

bharat bajaj IIT Delhi
10 years ago
The capital letters denote the position vectors of the points:
G = A + B + C/3
G1 = P + Q + R/3
3(G-G1) = A-P + B - Q + C - R
3GG1 = AP + BQ + CR

Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi